OpenStudy (anonymous):

solve each equation. state the number and type of roots. PROBLEM: x^(5)-81x=0 answer pwease....and explanation :D

OpenStudy (anonymous):

x(x^2+9)*(x^2-9) = 0 x=0, +/-3

OpenStudy (anonymous):

x^5 - 81x = x*(x^4 - 81) = x*(x^2 + 9)*(x^2-9) =x*(x^2+9)*(x+3)*(x-3) so as you can see from the factors, there are 3 real roots and two imaginary roots for a total of 5.

OpenStudy (anonymous):

AHHHH thanks you my friend, ur such a good help ! ^_~

OpenStudy (anonymous):

np & yw